OpenStudy (anonymous):

Use implicit diffentiation to find an equation of the line tangent to the given curve at the point (2,1) sin(xy-2)+x-2=(y-1)cos(x-2y) @perl :)

OpenStudy (anonymous):

Differentiation*

zepdrix (zepdrix):

Hey c: So where we stuck? Tried taking a derivative yet?

OpenStudy (anonymous):

Doing it right now, I got scared because of the wording of the problem but I think it's actually pretty simple. Hold on! :)

OpenStudy (anonymous):

Taking the deriv

OpenStudy (anonymous):

With a problem like this can I assume it's with respect to x or is that a nono?

zepdrix (zepdrix):

Yes, that's fine :) Take derivative with respect to x. We assume y is y(x), a function of x, even though it's not written explicitly.

OpenStudy (anonymous):

<3 I'll get back to ya if I get stuck!

zepdrix (zepdrix):

k np. boy that derivative is going to be a doozy lol especially on the right hand side

OpenStudy (anonymous):

cos(xy-2)*y'+1=(y'(cosx-2y)+(y-1)(-sinx-2)*y' Did I screw it up?

OpenStudy (anonymous):

Extra ( whoops

zepdrix (zepdrix):

Both sides look a tiny bit off.. let's start with the left side.

OpenStudy (anonymous):

Ohh I needed mult rule on left I think

OpenStudy (ribhu):

@tjb69812 may i help u

zepdrix (zepdrix):

\[\Large\rm \cos(xy-2)\color{royalblue}{(xy-2)'}+1\]

zepdrix (zepdrix):

Yah chain rule for this first term in the blue :)

OpenStudy (anonymous):

1(y)+x(y')

zepdrix (zepdrix):

sorry product rule* yes what you said :)

zepdrix (zepdrix):

mm looks good!

zepdrix (zepdrix):

\[\Large\rm \cos(xy-2)\color{orangered}{(y+xy')}+1\]

zepdrix (zepdrix):

On the right side:\[\Large\rm (y-1)\cos(x-2y)\]Product rule again, looks like you got that started correctly,\[\Large\rm \color{royalblue}{(y-1)'}\cos(x-2y)+(y-1)\color{royalblue}{(\cos(x-2y))'}\]And then I think you must made a small error on the end there

OpenStudy (anonymous):

y'*cos(x-2y)+(y-1)(-sin(x-2y)*[(1-2)*y'] ???

OpenStudy (anonymous):

Hahah it's late I may be goofing up again

zepdrix (zepdrix):

mmm that looks closer:\[\large\rm \color{orangered}{(y')}\cos(x-2y)+(y-1)\color{orangered}{(-\sin(x-2y))\color{royalblue}{(x-2y)'}}\] \[\large\rm \color{orangered}{(y')}\cos(x-2y)+(y-1)\color{orangered}{(-\sin(x-2y))\color{orangered}{(1-2y')}}\]

zepdrix (zepdrix):

That last y' should be inside the brackets, yah?

OpenStudy (anonymous):

Yessir you're correct on that

zepdrix (zepdrix):

So we've got something like this: \[\rm \cos(xy-2)(y+xy')+1=(y')\cos(x-2y)+(y-1)(-\sin(x-2y))(1-2y')\]

OpenStudy (anonymous):

Time to plug in?

OpenStudy (anonymous):

Or wait isolate y' right

zepdrix (zepdrix):

We COULD solve for y' and see that it's a function involving both x and y. So we'll think of y' as \(\Large\rm y'(x,y)\). BUT solving for y' is soooo much work here. Plugging in first will simplify things quite a bit I think.

OpenStudy (anonymous):

cos(2*1*-2)*(1+2*y')+1 = y' * cos(2-2*1)

OpenStudy (anonymous):

Doesn't the last term just = 0 that's what took me a second hahah

OpenStudy (anonymous):

Yea nvm it does

zepdrix (zepdrix):

\[\Large\rm \cos(2-2)(1+2y')=y' \cos(2-2)+(1-1)...\] Yah we can totally ignore that last guy :)

zepdrix (zepdrix):

\[\Large\rm \cos0(1+2y')=y' \cos0\]

zepdrix (zepdrix):

Shouldn't be too hard to solve from there, just a few algebra steps, yah?

OpenStudy (anonymous):

Safe to plug in 1 for cos0?

zepdrix (zepdrix):

Yup, that looks accurate!

zepdrix (zepdrix):

\[\Large\rm 1+2y'=y'\]

OpenStudy (anonymous):

I think we dropped a 1 somewhere or am I seeing things

OpenStudy (anonymous):

(1+2*y')+1 on the end

zepdrix (zepdrix):

On the left side? Ooo good call

zepdrix (zepdrix):

my beeeaaddd,\[\Large\rm \cos(2-2)(1+2y')+1=y' \cos(2-2)+(1-1)...\]\[\Large\rm \cos0(1+2y')+1=y'\cos0\]\[\Large\rm 1(1+2y')+1=y'\cdot 1\]\[\Large\rm 1+2y'+1=y'\]

OpenStudy (anonymous):

I'm getting y' = -2 after that yeah

OpenStudy (anonymous):

so y-1=-2(x-2) :D

zepdrix (zepdrix):

So y' gave you your slope? Good good good? Writing it in point-slope form? Yah that seems like the best idea here :) Since we're given a point. Yayyy team! \c:/ Good job!

OpenStudy (anonymous):

Thanks a ton dude!

zepdrix (zepdrix):

np